OpenStudy (adunb8):

xc = (y^2/x^2 -1)^1/2 how do you solve for y?

OpenStudy (anonymous):

You have to "cancel" the operation of square root - and then other actions. But first how does one neutralize the square root ? ....this is hint

OpenStudy (anonymous):

YOU RAISE TO THE SQUARE of course

OpenStudy (adunb8):

oh i see ! !! square both sides lol

OpenStudy (anonymous):

This is tricky : IT USUSALLY "GIVES BIRTH" to unexistent additional solutions

OpenStudy (adunb8):

???

OpenStudy (anonymous):

like x= 2 has only one TRUE solution, but if square both sides then x^2 = 4 has two, one of which is artificial and does not exist in the original equation

OpenStudy (adunb8):

i dont understand how my teacher got \[y = \pm x \sqrt{cx^2+1}\]

OpenStudy (anonymous):

He actually squared AND decided that BOTH solutions are legitimate.

OpenStudy (adunb8):

can you show me the steps?

OpenStudy (anonymous):

After squaring one obtainc y^2/x^2 - 1 = cx^2

OpenStudy (anonymous):

transfer 1 leftwise

OpenStudy (anonymous):

Multiply BOTH sdes by x^2

OpenStudy (adunb8):

wouldnt it be x^2c^2?

OpenStudy (anonymous):

Yes it must

OpenStudy (anonymous):

the way u wrote it the solution of ur teacher is a wee-bit missing the truth

OpenStudy (anonymous):

Hell-oo ?

OpenStudy (adunb8):

i still dont understand..

OpenStudy (adunb8):

because when i did it i got like something way off

OpenStudy (anonymous):

\[c^2x^{2} +1 = \frac{ y^2 }{ x^2 }\]

OpenStudy (anonymous):

Now multpl. both sides by x^2

OpenStudy (adunb8):

okay then that would be x^4c^2 + x^2

OpenStudy (anonymous):

One obtains \[x^2(c^2x^2 +1) = y^2 \]

OpenStudy (anonymous):

YOU DO NOT OPEN THE BRACKETS

OpenStudy (anonymous):

Extract the ROOT ! (like "RELEASE THE CRACKEN !")

OpenStudy (adunb8):

oh i see thank you! so you do sqrt to all of them then